Thomlin Company forecasts that total overhead for the current year will be $11,070,000 with 175,000 total machine hours. Year to date, the actual overhead is $7,860,000 and the actual machine hours are 81,000 hours. The predetermined overhead rate based on machine hours is

Round the factory overhead rate to the nearest dollar before multiplying by the number of hours.

a. $45 per machine hour b. $63 per machine hour c. $137 per machine hour d. $97 per machine hour

-----------------

Given the following information, determine the activity rate for setups.

Activity Total Activity-Base Usage Budgeted Activity Cost
Setups 9,600             $64,500            
Inspections 23,900             $138,600            
Assembly (dlh) 72,700             $405,000            

a. $5.80

b. $5.73

c. $5.57

d. $6.72

-----------------

Panamint Systems Corporation is estimating activity costs associated with producing disk drives, tapes drives, and wire drives. The indirect labor can be traced to four separate activity pools. The budgeted activity cost and activity base data by product are provided below.


Activity Cost

Activity Base
Procurement $398,700 Number of purchase orders
Scheduling 212,500 Number of production orders
Materials handling 433,900 Number of moves
Product development 720,500 Number of engineering changes
Production 1,431,300 Machine hours

Number of
Purchase
Orders

Number of
Production
Orders

Number
of
Moves

Number of
Engineering
Changes


Machine
Hours

Number
of
Units
Disk drives 3,990    270       1,500 15         1,800 2,300
Tape drives 2,500    205       690 7         10,000 4,500
Wire drives 12,700    830       4,300 23         10,300 2,100

The activity rate for the materials handling cost pool is

a. $66.86 per move

b. $64.76 per move

c. $20.78 per move

d. $162.84 per move

-------------------------

Botosan Factory has budgeted factory overhead for the year at $698,112, and budgeted direct labor hours for the year are 484,800. If the actual direct labor hours for the month of May are 441,200, the overhead allocated for May is

a. $540,029

b. $635,328

c. $654,388

d. $768,747

------------------------

Reynolds Manufacturers Inc. has estimated total factory overhead costs of $141,600 and expected direct labor hours of 11,800 for the current fiscal year. If job number 117 incurs 1,240 direct labor hours, Work in Process will be debited and Factory Overhead will be credited for

a. $14,880

b. $70,800

c. $141,600

d. $1,240
Answer #1

1) Overhead rate = 11070000/175000 = 63.26 per hour

So answer is b) $63 per hour

2) Setup rate = 64500/9600 = 6.72 per setup

So answer is d) $6.72

3) Material handling rate = 433900/6490 = 66.86 per move

So answer is a) $66.86 per move

4) Overhead rate = 698112/484800 = 1.44

Overhead allocated = 441200*1.44 = 635328

So answer is b) $635328

5) Overhead rate = 141600/11800 = 12

Overhead allocated = 1240*12 = 14880

So answer is a) $14880